OpenStudy (anonymous):

This one is challenging, at least for me, I know you guys are brainiacs..:>I don't think my answer is right though. here it is: Problem: r^2-6tw+2wr-3tr my answer: (r+2w)((r-3t)

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

thanks peeps

OpenStudy (anonymous):

yeah do the back calculation on your answer, you will get the question :)

OpenStudy (anonymous):

Factor by grouping: r^2-6tw+2wr-3tr = r^2+2wr -6tw - 3tr =r(r+2w) - 3t(r+2w) =(r+2w)(r-3t) =your answer!

OpenStudy (anonymous):

thanks fun. Thanks a bunch!!!!!!